how far will a car moving at 70km/h go in 30 minutes

Answer by TimothyLamb(4379)

---
s = d / t
d = s * t
---
d = 70(kph) * 0.5(hr)
d = 35 km
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php